I need help solving by using the quadratic formula
3x2-x=3

Answer by Edwin McCravy(20086)


I need help solving by using the quadratic formula 

              3x² - x = 3 

Get 0 on the right by adding -3 to both sides:

          3x² - x - 3 = 0

Use the quadratic formula, which gives 
the solution to

          ax² + bx + c = 0

The quadratic formula is
                  ______ 
            -b ± Öb²-4ac
        x = —————————————
                2a 

Compare 

           3x² - x - 3 = 0
to
           ax² + bx + c = 0

and we see that:  a = 3; b = -1; c = -3

                      ______________
             -(-1) ± Ö(-1)²-4(3)(-3)
        x = ————————————————————————
                     2(3) 
                  ____ 
             1 ± Ö1+36
        x = ————————————
                  6

                  __ 
             1 ± Ö37
        x = ———————————
                 6 

                      
Edwin